OpenStudy (anonymous):

What is the second step to solve the equation c divided by negative three plus eight equals two ?

OpenStudy (anonymous):

Multiply by negative three Add 8 Divide by negative three Subtract 8

OpenStudy (anonymous):

@Easyaspi314

OpenStudy (anonymous):

Does it look like this?

OpenStudy (anonymous):

\[\frac{ c }{ (-3+8) } = 2\]

OpenStudy (johnweldon1993):

Based on the text I'm going to say it's \[\frac{c}{-3} + 8 = 2\] Right?

OpenStudy (anonymous):

it looks like how @johnweldon1993 did it.

OpenStudy (anonymous):

@johnweldon1993 help please

OpenStudy (anonymous):

What did you do first?

OpenStudy (anonymous):

multiply by 3

OpenStudy (johnweldon1993):

No you want to isolate that fraction first...you need to do something with the 8 first right...?

OpenStudy (anonymous):

ohhh thats what i was confused about! so the second step would be to divide by 3!

OpenStudy (anonymous):

Not quite. Since you are already dividing by -3 in the fraction, you want to make sure you do the inverse on both sides of the equation.

OpenStudy (johnweldon1993):

So your first step would be subtract the 8 from both sides....so now you would have \[\large \frac{c}{-3} = -6\] what next?

OpenStudy (johnweldon1993):

*Hint...see how the variable is being DIVIDED by -3? well to eliminate that you would do the inverse of division...which is...?

OpenStudy (anonymous):

multiply

OpenStudy (johnweldon1993):

Correct so since this is the second step...your answer is...?

OpenStudy (anonymous):

so the answer would be multiplying by 3! Thank you

OpenStudy (johnweldon1993):

Correct!

OpenStudy (johnweldon1993):

Well 'negative' 3 but yes...

OpenStudy (anonymous):

Always watch your signs!